The centers of two masses are 5m apart. The masses exert a gravitational force of
2.4x10^-4 N on each other. Find the masses if one is three times the mass of the other.
1
Expert's answer
2021-10-20T16:39:52-0400

"F=G\frac{m\cdot3m}{x^2}\to m=\sqrt{\frac{Fx^2}{3G}}=\sqrt{\frac{2.4\cdot10^{-4}\cdot5^2}{3\cdot 6.67\cdot10^{-11}}}=5476\ (kg)"